2011-06-12 (UT)
==========

Program: CL4 (YSOs, PI: Kraus et al.)

Observers: Fabien Baron, Chris Farrington, FLUOR team

B1  B2  B3
----------
E1  W1  W2   (W1=REF)


(S1S2E1E2W1W2: 4-5-1-3-1-5)

CLIMB on beam 1,2,3

LT  - 20:40

*** E1W1W2, W1=ref.


Target HD142666, Cals: 140990 and 143616

start: weather clear, humidity 40-70%
r0 = 6-8 cm
scan length LONG

check star NIRO ALIGN on HD 141513

20:55 110/-1700
21:08 200/-1820      

Not sucessful finding W1W2 on calibrators, even if NIRO ALIGN on cal

Had to stop due to humidity now reaching 85%+ anyway

23:20 reopening, trying mwc275

      NIRO align on cal 164104
      -236/-1790
FLUOR doing VEGA, had to debias tip/tilt

23:30 -220/-1653
      500 /DEST (750 NONDEST not possible, saturated)
      recording HD 164104  (used previously) 
      counts: 370 /400/300
      V: 0.22 / 0.40 / 0.48
      check beam 1 shutter at the end (fluke)
23:38 -200/-1590      
      scrapping the previous cal, seems like a variable (flux way higher than others)

--

0:32 now trying v1295 Aql, cals HD188385 and 189509
0:50 found ?/-1780     
     testing modes on cal while waiting
     250 DEST ok
     500 NONDEST better 
1:24 now in delay -180/-1530
     recording fringes on HD188385 at 500/NONDEST
     counts: 60 / 40 / 30
     V: 0.50 / 0.52 / 0.35
1:54 -300/-1384
2:01 recording on v1295Aql, but saturating and not seeing fringes involving E1...
     piston seems too high to do anything
     

     trying MWC361 as backup
2:30 -1900/-2500
     calib HD 198236
     counts: 180 / 210 / 190
     V: 0.49 / 0.52 / 0.36
2:35 -1891/-2500
2:50 recording fringes on MWC361 (using env COG)
     beginning does not have much signal, then gets much better with all 3T
     counts: 105/120/90
     V: 0.12 / 0.20 / 0.16
2:59 -1773/-2400
     cal HD 197950
     counts: 75 / 90 / 75
     V: 0.35 / 0.35 / 0.35
3:19 MWC 361
     -1890/-2370
     counts: 120 / 140 / 95
     V: 0.10 / 0.23/ 0.11
3:32 cal HD 198236 (note: odd looking powerspectrum -- will need to investigate)
     -2000/-2480
     counts: 120 / 170 / 140
     V: 0.26 / 0.27 / 0.19 
NIRO ALIGN
     
3:51  cal HD 198236
      -2040 / -2420 
      counts: 170 / 190 /140
      V: 0.30 / 0.36 / 0.30
4:03  MWC 361
      -1916/-2360
      counts: 100 / 130/ 100
      V: 0.10 / 0.23 / 0.11
4:15  cal HD 197950
      -1920 / -2300
      counts: 80 / 100 / 80
      V: 0.37 / 0.33 / 0.37
4:30  MWC 361
      -1950/-2350
      counts: 110 / 130 / 90
      V: 0.11 / 0.25 / 0.12
4:45  cal HD 198236 
      -2090/-2440
      counts: 165/ 210 / 165
      V: 0.23 / 0.23 /0.24
5:00  MWC 361
      -2000/-2380
      counts: 120 / 130 / 100
      V: 0.08 / 0.27 / 0.10     

----
summary of observations:

Humidity levels rose during the first hours, preventing acquisition of early targets. 
After we reopened, not enough time was left to secure brackets on MWC275 or v1295Aql.
Due to high piston (but ok seeing), we had to settle for MWC 361 once again.

MWC 361 (5 brackets)